What is the equation of a line that passes through (7, -2) and has a slope of 3?
Answer by nerdybill(7384)

What is the equation of a line that passes through (7, -2) and has a slope of 3?
plug into "point-slope" form:
y - y1 = m(x - x1)
y - (-2) = 3(x - 7)
y + 2 = 3x - 21
y = 3x - 23 (this is what they're looking for)